### v3

# The T2K Collaboration First differential measurement of the single production cross section in neutrino neutral-current scattering K. Abe

††thanks: also at Université Paris-Saclay††thanks: also at Departament de Fisica de la Universitat Autonoma de Barcelona, Barcelona, Spain††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at Moscow Institute of Physics and Technology (MIPT), Moscow region, Russia and National Research Nuclear University ”MEPhI”, Moscow, Russia††thanks: also at IPSA-DRII, France††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at JINR, Dubna, Russia††thanks: also at the Graduate University of Science and Technology, Vietnam Academy of Science and Technology††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at Nambu Yoichiro Institute of Theoretical and Experimental Physics (NITEP)††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan††thanks: also at J-PARC, Tokai, Japan††thanks: also at Nambu Yoichiro Institute of Theoretical and Experimental Physics (NITEP)††thanks: affiliated member at Kavli IPMU (WPI), the University of Tokyo, Japan
The T2K Collaboration
First differential measurement of the single production cross section in neutrino neutral-current scattering
K. Abe
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
S. Abe
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
R. Akutsu
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
H. Alarakia-Charles
Lancaster University, Physics Department, Lancaster, United Kingdom
Y.I. Alj Hakim
University of Sheffield, School of Mathematical and Physical Sciences, Sheffield, United Kingdom
S. Alonso Monsalve
ETH Zurich, Institute for Particle Physics and Astrophysics, Zurich, Switzerland
L. Anthony
Imperial College London, Department of Physics, London, United Kingdom
S. Aoki
Kobe University, Kobe, Japan
K.A. Apte
Imperial College London, Department of Physics, London, United Kingdom
T. Arai
University of Tokyo, Department of Physics, Tokyo, Japan
T. Arihara
Tokyo Metropolitan University, Department of Physics, Tokyo, Japan
S. Arimoto
Kyoto University, Department of Physics, Kyoto, Japan
Y. Ashida
Tohoku University, Faculty of Science, Department of Physics, Miyagi, Japan
E.T. Atkin
Imperial College London, Department of Physics, London, United Kingdom
N. Babu
Louisiana State University, Department of Physics and Astronomy, Baton Rouge, Louisiana, U.S.A.
V. Baranov
Joint Institute for Nuclear Research, Dubna, Moscow Region, Russia
G.J. Barker
University of Warwick, Department of Physics, Coventry, United Kingdom
G. Barr
Oxford University, Department of Physics, Oxford, United Kingdom
D. Barrow
Oxford University, Department of Physics, Oxford, United Kingdom
P. Bates
University of Liverpool, Department of Physics, Liverpool, United Kingdom
L. Bathe-Peters
Oxford University, Department of Physics, Oxford, United Kingdom
M. Batkiewicz-Kwasniak
H. Niewodniczanski Institute of Nuclear Physics PAN, Cracow, Poland
N. Baudis
Oxford University, Department of Physics, Oxford, United Kingdom
V. Berardi
INFN Sezione di Bari and Università e Politecnico di Bari, Dipartimento Interuniversitario di Fisica, Bari, Italy
L. Berns
Tohoku University, Faculty of Science, Department of Physics, Miyagi, Japan
S. Bhattacharjee
Louisiana State University, Department of Physics and Astronomy, Baton Rouge, Louisiana, U.S.A.
A. Blanchet
CERN European Organization for Nuclear Research, CH-1211 Genéve 23, Switzerland
A. Blondel
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
S. Bolognesi
IRFU, CEA, Université Paris-Saclay, F-91191 Gif-sur-Yvette, France
S. Bordoni
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
S.B. Boyd
University of Warwick, Department of Physics, Coventry, United Kingdom
C. Bronner
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
A. Bubak
University of Silesia, Institute of Physics, Katowice, Poland
M. Buizza Avanzini
Ecole Polytechnique, IN2P3-CNRS, Laboratoire Leprince-Ringuet, Palaiseau, France
J.A. Caballero
Departamento de Física Atómica, Molecular y Nuclear, Universidad de Sevilla, 41080 Sevilla, Spain
F. Cadoux
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
N.F. Calabria
INFN Sezione di Bari and Università e Politecnico di Bari, Dipartimento Interuniversitario di Fisica, Bari, Italy
S. Cao
Institute For Interdisciplinary Research in Science and Education (IFIRSE), ICISE, Quy Nhon, Vietnam
S. Cap
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
D. Carabadjac
Ecole Polytechnique, IN2P3-CNRS, Laboratoire Leprince-Ringuet, Palaiseau, France
S.L. Cartwright
University of Sheffield, School of Mathematical and Physical Sciences, Sheffield, United Kingdom
M.P. Casado
Institut de Fisica d’Altes Energies (IFAE) - The Barcelona Institute of Science and Technology, Campus UAB, Bellaterra (Barcelona) Spain
M.G. Catanesi
INFN Sezione di Bari and Università e Politecnico di Bari, Dipartimento Interuniversitario di Fisica, Bari, Italy
J. Chakrani
Lawrence Berkeley National Laboratory, Berkeley, California, U.S.A.
A. Chalumeau
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
A. Chvirova
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
G. Collazuol
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
F. Cormier
TRIUMF, Vancouver, British Columbia, Canada
A.A.L. Craplet
Imperial College London, Department of Physics, London, United Kingdom
A. Cudd
University of Colorado at Boulder, Department of Physics, Boulder, Colorado, U.S.A.
D. D’ago
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
C. Dalmazzone
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
T. Daret
IRFU, CEA, Université Paris-Saclay, F-91191 Gif-sur-Yvette, France
P. Dasgupta
Eötvös Loránd University, Department of Atomic Physics, Budapest, Hungary
C. Davis
University of Pennsylvania, Department of Physics and Astronomy, Philadelphia, Pennsylvania, U.S.A.
Yu.I. Davydov
Joint Institute for Nuclear Research, Dubna, Moscow Region, Russia
G. De Rosa
INFN Sezione di Napoli and Università di Napoli, Dipartimento di Fisica, Napoli, Italy
T. Dealtry
Lancaster University, Physics Department, Lancaster, United Kingdom
C. Densham
STFC, Rutherford Appleton Laboratory, Harwell Oxford, and Daresbury Laboratory, Warrington, United Kingdom
A. Dergacheva
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
R. Dharmapal Banerjee
Wroclaw University, Faculty of Physics and Astronomy, Wroclaw, Poland
F. Di Lodovico
King’s College London, Department of Physics, Strand, London WC2R 2LS, United Kingdom
G. Diaz Lopez
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
S. Dolan
CERN European Organization for Nuclear Research, CH-1211 Genéve 23, Switzerland
D. Douqa
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
T.A. Doyle
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
O. Drapier
Ecole Polytechnique, IN2P3-CNRS, Laboratoire Leprince-Ringuet, Palaiseau, France
K.E. Duffy
Oxford University, Department of Physics, Oxford, United Kingdom
J. Dumarchez
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
P. Dunne
Imperial College London, Department of Physics, London, United Kingdom
K. Dygnarowicz
Warsaw University of Technology, Institute of Radioelectronics and Multimedia Technology, Warsaw, Poland
A. Eguchi
University of Tokyo, Department of Physics, Tokyo, Japan
J. Elias
University of Rochester, Department of Physics and Astronomy, Rochester, New York, U.S.A.
S. Emery-Schrenk
IRFU, CEA, Université Paris-Saclay, F-91191 Gif-sur-Yvette, France
G. Erofeev
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
A. Ershova
Ecole Polytechnique, IN2P3-CNRS, Laboratoire Leprince-Ringuet, Palaiseau, France
G. Eurin
IRFU, CEA, Université Paris-Saclay, F-91191 Gif-sur-Yvette, France
D. Fedorova
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
S. Fedotov
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
M. Feltre
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
L. Feng
Kyoto University, Department of Physics, Kyoto, Japan
D. Ferlewicz
University of Tokyo, Department of Physics, Tokyo, Japan
A.J. Finch
Lancaster University, Physics Department, Lancaster, United Kingdom
M.D. Fitton
STFC, Rutherford Appleton Laboratory, Harwell Oxford, and Daresbury Laboratory, Warrington, United Kingdom
C. Forza
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
M. Friend
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
Y. Fujii
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
Y. Fukuda
Miyagi University of Education, Department of Physics, Sendai, Japan
Y. Furui
Tokyo Metropolitan University, Department of Physics, Tokyo, Japan
J. García-Marcos
Ghent University, Department of Physics and Astronomy, Proeftuinstraat 86, B-9000 Gent, Belgium
A.C. Germer
University of Pennsylvania, Department of Physics and Astronomy, Philadelphia, Pennsylvania, U.S.A.
L. Giannessi
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
C. Giganti
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
V. Glagolev
Joint Institute for Nuclear Research, Dubna, Moscow Region, Russia
M. Gonin
ILANCE, CNRS – University of Tokyo International Research Laboratory, Kashiwa, Chiba 277-8582, Japan
R. González Jiménez
Departamento de Física Atómica, Molecular y Nuclear, Universidad de Sevilla, 41080 Sevilla, Spain
J. González Rosa
Departamento de Física Atómica, Molecular y Nuclear, Universidad de Sevilla, 41080 Sevilla, Spain
E.A.G. Goodman
University of Glasgow, School of Physics and Astronomy, Glasgow, United Kingdom
K. Gorshanov
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
P. Govindaraj
University of Warsaw, Faculty of Physics, Warsaw, Poland
M. Grassi
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
M. Guigue
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
F.Y. Guo
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
D.R. Hadley
University of Warwick, Department of Physics, Coventry, United Kingdom
S. Han
Kyoto University, Department of Physics, Kyoto, Japan
University of Tokyo, Institute for Cosmic Ray Research, Research Center for Cosmic Neutrinos, Kashiwa, Japan
D.A. Harris
York University, Department of Physics and Astronomy, Toronto, Ontario, Canada
R.J. Harris
Lancaster University, Physics Department, Lancaster, United Kingdom
STFC, Rutherford Appleton Laboratory, Harwell Oxford, and Daresbury Laboratory, Warrington, United Kingdom
T. Hasegawa
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
C.M. Hasnip
CERN European Organization for Nuclear Research, CH-1211 Genéve 23, Switzerland
S. Hassani
IRFU, CEA, Université Paris-Saclay, F-91191 Gif-sur-Yvette, France
N.C. Hastings
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
Y. Hayato
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
I. Heitkamp
Tohoku University, Faculty of Science, Department of Physics, Miyagi, Japan
D. Henaff
IRFU, CEA, Université Paris-Saclay, F-91191 Gif-sur-Yvette, France
Y. Hino
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
J. Holeczek
University of Silesia, Institute of Physics, Katowice, Poland
A. Holin
STFC, Rutherford Appleton Laboratory, Harwell Oxford, and Daresbury Laboratory, Warrington, United Kingdom
T. Holvey
Oxford University, Department of Physics, Oxford, United Kingdom
N.T. Hong Van
International Centre of Physics, Institute of Physics (IOP), Vietnam Academy of Science and Technology (VAST), 10 Dao Tan, Ba Dinh, Hanoi, Vietnam
T. Honjo
Osaka Metropolitan University, Department of Physics, Osaka, Japan
M.C.F. Hooft
Ghent University, Department of Physics and Astronomy, Proeftuinstraat 86, B-9000 Gent, Belgium
K. Hosokawa
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
J. Hu
Kyoto University, Department of Physics, Kyoto, Japan
A.K. Ichikawa
Tohoku University, Faculty of Science, Department of Physics, Miyagi, Japan
K. Ieki
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
M. Ikeda
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
T. Ishida
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
M. Ishitsuka
Tokyo University of Science, Faculty of Science and Technology, Department of Physics, Noda, Chiba, Japan
A. Izmaylov
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
N. Jachowicz
Ghent University, Department of Physics and Astronomy, Proeftuinstraat 86, B-9000 Gent, Belgium
S.J. Jenkins
University of Liverpool, Department of Physics, Liverpool, United Kingdom
C. Jesús-Valls
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
M. Jia
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
J.J. Jiang
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
J.Y. Ji
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
T.P. Jones
Lancaster University, Physics Department, Lancaster, United Kingdom
P. Jonsson
Imperial College London, Department of Physics, London, United Kingdom
S. Joshi
IRFU, CEA, Université Paris-Saclay, F-91191 Gif-sur-Yvette, France
M. Kabirnezhad
Imperial College London, Department of Physics, London, United Kingdom
A.C. Kaboth
Royal Holloway University of London, Department of Physics, Egham, Surrey, United Kingdom
STFC, Rutherford Appleton Laboratory, Harwell Oxford, and Daresbury Laboratory, Warrington, United Kingdom
H. Kakuno
Tokyo Metropolitan University, Department of Physics, Tokyo, Japan
J. Kameda
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
S. Karpova
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
V.S. Kasturi
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
Y. Kataoka
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
T. Katori
King’s College London, Department of Physics, Strand, London WC2R 2LS, United Kingdom
Y. Kawamura
Osaka Metropolitan University, Department of Physics, Osaka, Japan
M. Kawaue
Kyoto University, Department of Physics, Kyoto, Japan
E. Kearns
Boston University, Department of Physics, Boston, Massachusetts, U.S.A.
M. Khabibullin
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
A. Khotjantsev
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
T. Kikawa
Kyoto University, Department of Physics, Kyoto, Japan
S. King
King’s College London, Department of Physics, Strand, London WC2R 2LS, United Kingdom
V. Kiseeva
Joint Institute for Nuclear Research, Dubna, Moscow Region, Russia
J. Kisiel
University of Silesia, Institute of Physics, Katowice, Poland
A. Klustová
Imperial College London, Department of Physics, London, United Kingdom
L. Kneale
University of Sheffield, School of Mathematical and Physical Sciences, Sheffield, United Kingdom
H. Kobayashi
University of Tokyo, Department of Physics, Tokyo, Japan
L. Koch
Institut für Physik, Johannes Gutenberg-Universität Mainz, Staudingerweg 7, 55128 Mainz, Germany
S. Kodama
University of Tokyo, Department of Physics, Tokyo, Japan
M. Kolupanova
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
A. Konaka
TRIUMF, Vancouver, British Columbia, Canada
L.L. Kormos
Lancaster University, Physics Department, Lancaster, United Kingdom
Y. Koshio
Okayama University, Department of Physics, Okayama, Japan
K. Kowalik
National Centre for Nuclear Research, Warsaw, Poland
Y. Kudenko
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
Y. Kudo
Yokohama National University, Department of Physics, Yokohama, Japan
A. Kumar Jha
Ghent University, Department of Physics and Astronomy, Proeftuinstraat 86, B-9000 Gent, Belgium
R. Kurjata
Warsaw University of Technology, Institute of Radioelectronics and Multimedia Technology, Warsaw, Poland
V. Kurochka
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
T. Kutter
Louisiana State University, Department of Physics and Astronomy, Baton Rouge, Louisiana, U.S.A.
L. Labarga
University Autonoma Madrid, Department of Theoretical Physics, 28049 Madrid, Spain
M. Lachat
University of Rochester, Department of Physics and Astronomy, Rochester, New York, U.S.A.
K. Lachner
University of Warwick, Department of Physics, Coventry, United Kingdom
J. Lagoda
National Centre for Nuclear Research, Warsaw, Poland
S.M. Lakshmi
University of Silesia, Institute of Physics, Katowice, Poland
M. Lamers James
University of Warwick, Department of Physics, Coventry, United Kingdom
A. Langella
INFN Sezione di Napoli and Università di Napoli, Dipartimento di Fisica, Napoli, Italy
D.H. Langridge
Royal Holloway University of London, Department of Physics, Egham, Surrey, United Kingdom
J.-F. Laporte
IRFU, CEA, Université Paris-Saclay, F-91191 Gif-sur-Yvette, France
D. Last
University of Rochester, Department of Physics and Astronomy, Rochester, New York, U.S.A.
N. Latham
King’s College London, Department of Physics, Strand, London WC2R 2LS, United Kingdom
M. Laveder
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
M. Lawe
Lancaster University, Physics Department, Lancaster, United Kingdom
D. Leon Silverio
South Dakota School of Mines and Technology, 501 East Saint Joseph Street, Rapid City, SD 57701, United States
S. Levorato
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
S.V. Lewis
King’s College London, Department of Physics, Strand, London WC2R 2LS, United Kingdom
B. Li
ETH Zurich, Institute for Particle Physics and Astrophysics, Zurich, Switzerland
C. Lin
Imperial College London, Department of Physics, London, United Kingdom
R.P. Litchfield
University of Glasgow, School of Physics and Astronomy, Glasgow, United Kingdom
S.L. Liu
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
W. Li
Oxford University, Department of Physics, Oxford, United Kingdom
A. Longhin
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
A. Lopez Moreno
King’s College London, Department of Physics, Strand, London WC2R 2LS, United Kingdom
L. Ludovici
INFN Sezione di Roma and Università di Roma “La Sapienza”, Roma, Italy
X. Lu
University of Warwick, Department of Physics, Coventry, United Kingdom
T. Lux
Institut de Fisica d’Altes Energies (IFAE) - The Barcelona Institute of Science and Technology, Campus UAB, Bellaterra (Barcelona) Spain
L.N. Machado
University of Glasgow, School of Physics and Astronomy, Glasgow, United Kingdom
L. Magaletti
INFN Sezione di Bari and Università e Politecnico di Bari, Dipartimento Interuniversitario di Fisica, Bari, Italy
K. Mahn
Michigan State University, Department of Physics and Astronomy, East Lansing, Michigan, U.S.A.
K.K. Mahtani
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
M. Mandal
National Centre for Nuclear Research, Warsaw, Poland
S. Manly
University of Rochester, Department of Physics and Astronomy, Rochester, New York, U.S.A.
A.D. Marino
University of Colorado at Boulder, Department of Physics, Boulder, Colorado, U.S.A.
D.G.R. Martin
Imperial College London, Department of Physics, London, United Kingdom
D.A. Martinez Caicedo
South Dakota School of Mines and Technology, 501 East Saint Joseph Street, Rapid City, SD 57701, United States
L. Martinez
Institut de Fisica d’Altes Energies (IFAE) - The Barcelona Institute of Science and Technology, Campus UAB, Bellaterra (Barcelona) Spain
M. Martini
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
T. Matsubara
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
R. Matsumoto
Institute of Science Tokyo, Department of Physics, Tokyo
V. Matveev
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
C. Mauger
University of Pennsylvania, Department of Physics and Astronomy, Philadelphia, Pennsylvania, U.S.A.
K. Mavrokoridis
University of Liverpool, Department of Physics, Liverpool, United Kingdom
N. McCauley
University of Liverpool, Department of Physics, Liverpool, United Kingdom
K.S. McFarland
University of Rochester, Department of Physics and Astronomy, Rochester, New York, U.S.A.
J. McKean
Imperial College London, Department of Physics, London, United Kingdom
A. Mefodiev
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
G.D. Megias
Departamento de Física Atómica, Molecular y Nuclear, Universidad de Sevilla, 41080 Sevilla, Spain
L. Mellet
Michigan State University, Department of Physics and Astronomy, East Lansing, Michigan, U.S.A.
M. Mezzetto
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
S. Miki
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
V. Mikola
University of Glasgow, School of Physics and Astronomy, Glasgow, United Kingdom
E.W. Miller
Institut de Fisica d’Altes Energies (IFAE) - The Barcelona Institute of Science and Technology, Campus UAB, Bellaterra (Barcelona) Spain
A. Minamino
Yokohama National University, Department of Physics, Yokohama, Japan
O. Mineev
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
S. Mine
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
University of California, Irvine, Department of Physics and Astronomy, Irvine, California, U.S.A.
J. Mirabito
Boston University, Department of Physics, Boston, Massachusetts, U.S.A.
M. Miura
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
S. Moriyama
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
S. Moriyama
Yokohama National University, Department of Physics, Yokohama, Japan
P. Morrison
University of Glasgow, School of Physics and Astronomy, Glasgow, United Kingdom
Th.A. Mueller
Ecole Polytechnique, IN2P3-CNRS, Laboratoire Leprince-Ringuet, Palaiseau, France
D. Munford
University of Houston, Department of Physics, Houston, Texas, U.S.A.
A. Muñoz
Ecole Polytechnique, IN2P3-CNRS, Laboratoire Leprince-Ringuet, Palaiseau, France
ILANCE, CNRS – University of Tokyo International Research Laboratory, Kashiwa, Chiba 277-8582, Japan
L. Munteanu
CERN European Organization for Nuclear Research, CH-1211 Genéve 23, Switzerland
Y. Nagai
Eötvös Loránd University, Department of Atomic Physics, Budapest, Hungary
T. Nakadaira
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
K. Nakagiri
University of Tokyo, Department of Physics, Tokyo, Japan
M. Nakahata
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
Y. Nakajima
University of Tokyo, Department of Physics, Tokyo, Japan
K.D. Nakamura
Tohoku University, Faculty of Science, Department of Physics, Miyagi, Japan
Y. Nakano
University of Toyama, Department of Physics, Toyama, Japan
S. Nakayama
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
T. Nakaya
Kyoto University, Department of Physics, Kyoto, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
K. Nakayoshi
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
C.E.R. Naseby
Imperial College London, Department of Physics, London, United Kingdom
D.T. Nguyen
VNU University of Science, Vietnam National University, Hanoi, Vietnam
V.Q. Nguyen
Ecole Polytechnique, IN2P3-CNRS, Laboratoire Leprince-Ringuet, Palaiseau, France
K. Niewczas
Ghent University, Department of Physics and Astronomy, Proeftuinstraat 86, B-9000 Gent, Belgium
S. Nishimori
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
Y. Nishimura
Keio University, Department of Physics, Kanagawa, Japan
Y. Noguchi
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
T. Nosek
National Centre for Nuclear Research, Warsaw, Poland
F. Nova
STFC, Rutherford Appleton Laboratory, Harwell Oxford, and Daresbury Laboratory, Warrington, United Kingdom
J.C. Nugent
Imperial College London, Department of Physics, London, United Kingdom
H.M. O’Keeffe
Lancaster University, Physics Department, Lancaster, United Kingdom
L. O’Sullivan
Institut für Physik, Johannes Gutenberg-Universität Mainz, Staudingerweg 7, 55128 Mainz, Germany
R. Okazaki
Keio University, Department of Physics, Kanagawa, Japan
W. Okinaga
University of Tokyo, Department of Physics, Tokyo, Japan
K. Okumura
University of Tokyo, Institute for Cosmic Ray Research, Research Center for Cosmic Neutrinos, Kashiwa, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
T. Okusawa
Osaka Metropolitan University, Department of Physics, Osaka, Japan
N. Onda
Kyoto University, Department of Physics, Kyoto, Japan
N. Ospina
INFN Sezione di Bari and Università e Politecnico di Bari, Dipartimento Interuniversitario di Fisica, Bari, Italy
L. Osu
Ecole Polytechnique, IN2P3-CNRS, Laboratoire Leprince-Ringuet, Palaiseau, France
Y. Oyama
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
V. Paolone
University of Pittsburgh, Department of Physics and Astronomy, Pittsburgh, Pennsylvania, U.S.A.
J. Pasternak
Imperial College London, Department of Physics, London, United Kingdom
M. Pfaff
Imperial College London, Department of Physics, London, United Kingdom
L. Pickering
STFC, Rutherford Appleton Laboratory, Harwell Oxford, and Daresbury Laboratory, Warrington, United Kingdom
B. Popov
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
A.J. Portocarrero Yrey
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
M. Posiadala-Zezula
University of Warsaw, Faculty of Physics, Warsaw, Poland
Y.S. Prabhu
University of Warsaw, Faculty of Physics, Warsaw, Poland
H. Prasad
Wroclaw University, Faculty of Physics and Astronomy, Wroclaw, Poland
F. Pupilli
INFN Sezione di Padova and Università di Padova, Dipartimento di Fisica, Padova, Italy
B. Quilain
ILANCE, CNRS – University of Tokyo International Research Laboratory, Kashiwa, Chiba 277-8582, Japan
Ecole Polytechnique, IN2P3-CNRS, Laboratoire Leprince-Ringuet, Palaiseau, France
P.T. Quyen
Institute For Interdisciplinary Research in Science and Education (IFIRSE), ICISE, Quy Nhon, Vietnam
E. Radicioni
INFN Sezione di Bari and Università e Politecnico di Bari, Dipartimento Interuniversitario di Fisica, Bari, Italy
B. Radics
York University, Department of Physics and Astronomy, Toronto, Ontario, Canada
M.A. Ramirez
University of Pennsylvania, Department of Physics and Astronomy, Philadelphia, Pennsylvania, U.S.A.
R. Ramsden
King’s College London, Department of Physics, Strand, London WC2R 2LS, United Kingdom
P.N. Ratoff
Lancaster University, Physics Department, Lancaster, United Kingdom
M. Reh
University of Colorado at Boulder, Department of Physics, Boulder, Colorado, U.S.A.
G. Reina
Institut für Physik, Johannes Gutenberg-Universität Mainz, Staudingerweg 7, 55128 Mainz, Germany
C. Riccio
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
D.W. Riley
University of Glasgow, School of Physics and Astronomy, Glasgow, United Kingdom
E. Rondio
National Centre for Nuclear Research, Warsaw, Poland
S. Roth
RWTH Aachen University, III. Physikalisches Institut, Aachen, Germany
N. Roy
York University, Department of Physics and Astronomy, Toronto, Ontario, Canada
A. Rubbia
ETH Zurich, Institute for Particle Physics and Astrophysics, Zurich, Switzerland
L. Russo
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
A. Rychter
Warsaw University of Technology, Institute of Radioelectronics and Multimedia Technology, Warsaw, Poland
W. Saenz
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
K. Sakashita
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
F. Sánchez
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
E.M. Sandford
University of Liverpool, Department of Physics, Liverpool, United Kingdom
Y. Sato
Tokyo University of Science, Faculty of Science and Technology, Department of Physics, Noda, Chiba, Japan
T. Schefke
Louisiana State University, Department of Physics and Astronomy, Baton Rouge, Louisiana, U.S.A.
C.M. Schloesser
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
K. Scholberg
Duke University, Department of Physics, Durham, North Carolina, U.S.A.
M. Scott
Imperial College London, Department of Physics, London, United Kingdom
Y. Seiya
Osaka Metropolitan University, Department of Physics, Osaka, Japan
T. Sekiguchi
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
H. Sekiya
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
T. Sekiya
Tokyo Metropolitan University, Department of Physics, Tokyo, Japan
D. Seppala
Michigan State University, Department of Physics and Astronomy, East Lansing, Michigan, U.S.A.
D. Sgalaberna
ETH Zurich, Institute for Particle Physics and Astrophysics, Zurich, Switzerland
A. Shaikhiev
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
M. Shiozawa
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
Y. Shiraishi
Okayama University, Department of Physics, Okayama, Japan
A. Shvartsman
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
N. Skrobova
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
K. Skwarczynski
Royal Holloway University of London, Department of Physics, Egham, Surrey, United Kingdom
D. Smyczek
RWTH Aachen University, III. Physikalisches Institut, Aachen, Germany
M. Smy
University of California, Irvine, Department of Physics and Astronomy, Irvine, California, U.S.A.
J.T. Sobczyk
Wroclaw University, Faculty of Physics and Astronomy, Wroclaw, Poland
H. Sobel
University of California, Irvine, Department of Physics and Astronomy, Irvine, California, U.S.A.
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
F.J.P. Soler
University of Glasgow, School of Physics and Astronomy, Glasgow, United Kingdom
A.J. Speers
Lancaster University, Physics Department, Lancaster, United Kingdom
R. Spina
INFN Sezione di Bari and Università e Politecnico di Bari, Dipartimento Interuniversitario di Fisica, Bari, Italy
A. Srivastava
Institut für Physik, Johannes Gutenberg-Universität Mainz, Staudingerweg 7, 55128 Mainz, Germany
P. Stowell
University of Sheffield, School of Mathematical and Physical Sciences, Sheffield, United Kingdom
Y. Stroke
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
I.A. Suslov
Joint Institute for Nuclear Research, Dubna, Moscow Region, Russia
A. Suzuki
Kobe University, Kobe, Japan
S.Y. Suzuki
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
M. Tada
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
S. Tairafune
Tohoku University, Faculty of Science, Department of Physics, Miyagi, Japan
A. Takeda
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
Y. Takeuchi
Kobe University, Kobe, Japan
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
H.K. Tanaka
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
H. Tanigawa
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
V.V. Tereshchenko
Joint Institute for Nuclear Research, Dubna, Moscow Region, Russia
N. Thamm
RWTH Aachen University, III. Physikalisches Institut, Aachen, Germany
N. Tran
Kyoto University, Department of Physics, Kyoto, Japan
T. Tsukamoto
High Energy Accelerator Research Organization (KEK), Tsukuba, Ibaraki, Japan
M. Tzanov
Louisiana State University, Department of Physics and Astronomy, Baton Rouge, Louisiana, U.S.A.
Y. Uchida
Imperial College London, Department of Physics, London, United Kingdom
M. Vagins
Kavli Institute for the Physics and Mathematics of the Universe (WPI), The University of Tokyo Institutes for Advanced Study, University of Tokyo, Kashiwa, Chiba, Japan
University of California, Irvine, Department of Physics and Astronomy, Irvine, California, U.S.A.
M. Varghese
Institut de Fisica d’Altes Energies (IFAE) - The Barcelona Institute of Science and Technology, Campus UAB, Bellaterra (Barcelona) Spain
I. Vasilyev
Joint Institute for Nuclear Research, Dubna, Moscow Region, Russia
G. Vasseur
IRFU, CEA, Université Paris-Saclay, F-91191 Gif-sur-Yvette, France
E. Villa
CERN European Organization for Nuclear Research, CH-1211 Genéve 23, Switzerland
University of Geneva, Section de Physique, DPNC, Geneva, Switzerland
U. Virginet
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
T. Vladisavljevic
STFC, Rutherford Appleton Laboratory, Harwell Oxford, and Daresbury Laboratory, Warrington, United Kingdom
T. Wachala
H. Niewodniczanski Institute of Nuclear Physics PAN, Cracow, Poland
D. Wakabayashi
Tohoku University, Faculty of Science, Department of Physics, Miyagi, Japan
H.T. Wallace
University of Sheffield, School of Mathematical and Physical Sciences, Sheffield, United Kingdom
J.G. Walsh
Michigan State University, Department of Physics and Astronomy, East Lansing, Michigan, U.S.A.
L. Wan
Boston University, Department of Physics, Boston, Massachusetts, U.S.A.
D. Wark
STFC, Rutherford Appleton Laboratory, Harwell Oxford, and Daresbury Laboratory, Warrington, United Kingdom
Oxford University, Department of Physics, Oxford, United Kingdom
M.O. Wascko
Oxford University, Department of Physics, Oxford, United Kingdom
STFC, Rutherford Appleton Laboratory, Harwell Oxford, and Daresbury Laboratory, Warrington, United Kingdom
A. Weber
Institut für Physik, Johannes Gutenberg-Universität Mainz, Staudingerweg 7, 55128 Mainz, Germany
R. Wendell
Kyoto University, Department of Physics, Kyoto, Japan
M.J. Wilking
School of Physics and Astronomy, University of Minnesota, Minneapolis, Minnesota, U.S.A.
C. Wilkinson
Lawrence Berkeley National Laboratory, Berkeley, California, U.S.A.
J.R. Wilson
King’s College London, Department of Physics, Strand, London WC2R 2LS, United Kingdom
K. Wood
Lawrence Berkeley National Laboratory, Berkeley, California, U.S.A.
C. Wret
Imperial College London, Department of Physics, London, United Kingdom
J. Xia
SLAC National Accelerator Laboratory, Stanford University, Menlo Park, California, U.S.A.
K. Yamamoto
Osaka Metropolitan University, Department of Physics, Osaka, Japan
T. Yamamoto
Osaka Metropolitan University, Department of Physics, Osaka, Japan
Y. Yang
Oxford University, Department of Physics, Oxford, United Kingdom
T. Yano
University of Tokyo, Institute for Cosmic Ray Research, Kamioka Observatory, Kamioka, Japan
N. Yershov
Institute for Nuclear Research of the Russian Academy of Sciences, Moscow, Russia
U. Yevarouskaya
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
M. Yokoyama
University of Tokyo, Department of Physics, Tokyo, Japan
Y. Yoshimoto
University of Tokyo, Department of Physics, Tokyo, Japan
N. Yoshimura
Kyoto University, Department of Physics, Kyoto, Japan
R. Zaki
York University, Department of Physics and Astronomy, Toronto, Ontario, Canada
A. Zalewska
H. Niewodniczanski Institute of Nuclear Physics PAN, Cracow, Poland
J. Zalipska
National Centre for Nuclear Research, Warsaw, Poland
G. Zarnecki
H. Niewodniczanski Institute of Nuclear Physics PAN, Cracow, Poland
J. Zhang
TRIUMF, Vancouver, British Columbia, Canada
University of British Columbia, Department of Physics and Astronomy, Vancouver, British Columbia, Canada
X.Y. Zhao
ETH Zurich, Institute for Particle Physics and Astrophysics, Zurich, Switzerland
H. Zheng
State University of New York at Stony Brook, Department of Physics and Astronomy, Stony Brook, New York, U.S.A.
H. Zhong
Kobe University, Kobe, Japan
T. Zhu
Imperial College London, Department of Physics, London, United Kingdom
M. Ziembicki
Warsaw University of Technology, Institute of Radioelectronics and Multimedia Technology, Warsaw, Poland
E.D. Zimmerman
University of Colorado at Boulder, Department of Physics, Boulder, Colorado, U.S.A.
M. Zito
Sorbonne Université, CNRS/IN2P3, Laboratoire de Physique Nucléaire et de Hautes Energies (LPNHE), Paris, France
S. Zsoldos
King’s College London, Department of Physics, Strand, London WC2R 2LS, United Kingdom
Abstract
Since its first observation in the 1970s, neutrino-induced neutral-current single positive pion production (NC1) has remained an elusive and poorly understood interaction channel. This process is a significant background in neutrino oscillation experiments and studying it further is critical for the physics program of next-generation accelerator-based neutrino oscillation experiments. In this Letter we present the first double-differential cross-section measurement of NC1 interactions using data from the ND280 detector of the T2K experiment collected in -beam mode. The measured flux-averaged integrated cross-section is . We compare the results on a hydrocarbon target to the predictions of several neutrino interaction generators and final-state interaction models. While model predictions agree with the differential results, the data shows a weak preference for a cross-section normalization approximately 30% higher than predicted by most models studied in this Letter.
Introduction. Neutrino physics plays a central role in modern high-energy physics. Following the discovery of neutrino oscillations Fukuda et al. (1998); Ahmad et al. (2002) several experiments are ongoing and planned with the aim of precisely characterizing the phenomenology of neutrino oscillations. Neutrino mixing is characterized by three-flavor mixing, depending on two mass splittings (, ) and described by the Pontecorvo-Maki-Nakagawa-Sakata (PMNS) matrix, consisting of three mixing angles (, , ) and one Charge-Parity (CP) violating phase (. Several parameters are already known with a few percent precision Esteban et al. (2024), but major unknowns remain: the sign of , known as the neutrino mass ordering; whether CP symmetry is violated or not, which is primarily determined by the value of ; and, the octant of , determining if is smaller than, larger than or equal to .
For existing and planned neutrino beam experiments, the experimental sensitivity to the unknown neutrino oscillation parameters becomes maximal for neutrino energies ranging from hundreds of MeV to a few GeV. In this energy range, multiple neutrino interaction mechanisms are possible, and poorly understood nuclear effects contribute significant systematic uncertainties Alvarez-Ruso et al. (2018); Andreopoulos et al. (2019); Aliaga et al. (2020); Abe et al. (2023a). Over the past two decades, more than twenty measurements have investigated neutrino single production. Of these, more than ten have studied neutral-current (NC) interactions Abratenko et al. (2024a); Acciarri et al. (2017); Nakayama et al. (2005); Abratenko et al. (2023); Wolcott et al. (2016); Aguilar-Arevalo et al. (2010, 2008); Adamson et al. (2016); Kullenberg et al. (2009); Acero et al. (2020); Kurimoto et al. (2010a, b); Abe et al. (2018a), with the remainder studying charged-current (CC) interactions Abratenko et al. (2024b); Mariani et al. (2011); Adams et al. (2019); Coplowe et al. (2020); McGivern et al. (2016); Le et al. (2015); Altinok et al. (2017); Aguilar-Arevalo et al. (2011a); Acero et al. (2023). Similarly, more than twenty analyses have examined single production, though exclusively through CC interactions Ramírez et al. (2023); Acciarri et al. (2018, 2014); Rodriguez et al. (2008); Hasegawa et al. (2005); Bercellie et al. (2023); Eberly et al. (2015); McGivern et al. (2016); Higuera et al. (2014); Le et al. (2019); Mislivec et al. (2018); Aguilar-Arevalo et al. (2011b, 2009); Hiraide et al. (2008); Abe et al. (2023b, 2021, 2017, 2016, 2020); Oshima et al. (2022); Astier et al. (2001). However, measurements of neutrino NC single production (NC1), primarily generated by interactions, have been notably absent from the literature.
In this Letter, we address this long-standing absence of results for this channel by presenting the first differential cross-section measurement of NC1 interactions, reported as a function of the pion momentum and angle.
Figure 1: Expected event rate, in arbitrary units, for CC0 including and Cherenkov thresholds, (T2K/Hyper-K) and CC-inclusive (DUNE) muon-neutrino interactions at the Hyper-K (top) and DUNE (bottom) experiments, shown as a function of true neutrino energy. Also overlaid is the expected rate of the NC1 background in each case. The oscillation parameters used are from NuFit5.0 Esteban et al. (2020), and use GENIEv3 Andreopoulos et al. (2010, 2015) truth information to predict the interaction rates using T2K Abe et al. (2013) and DUNE Abi et al. (2020a, b) neutrino fluxes.
Existing Measurements. The Gargamelle Krenz et al. (1978) and ANL Derrick et al. (1980) bubble chamber experiments measured NC1 interactions over four decades ago. At the time, their focus was demonstrating consistency with the newly proposed Standard Model of particle physics, confirming the existence of these interactions. One measurement of the NC1 cross section exists in the scientific literatureiiiTo the best knowledge of the authors, details on the analysis methodology have been lost., a re-analysis of the Gargamelle reported event rates, collected at an average neutrino energy of 2 GeV Zeller (2003).
NC1 in LBE. Fig. 1 shows the predicted oscillated muon-neutrino interaction rate as a function of neutrino energy for two future long-baseline neutrino experiments (LBE): Hyper-Kamiokande Abe et al. (2018b), with an unoscillated neutrino energy peak at 0.6 GeV and a water Cherenkov far detector, which focuses on CC0 events; and DUNE Abi et al. (2020a), with a peak unoscillated neutrino energy of 2.5 GeV and a liquid argon far detector, using all CC events. The neutrino energy peak is deliberately aligned near the energy where maximum disappearance of the muon flavor is expected for the experiment’s baseline, creating a characteristic “dip” in the observed energy distribution. The depth of this dip and its position are used to determine the value of the and parameters. Overlaid is NC1 predicted event rate, which is a hard to reject background as pions are frequently mis-identified as muons in both water Cherenkov detectors and liquid argon detectors. Due to oscillations, the CC event rate associated to muon-neutrinos is strongly reduced near the dip, however, all neutrino flavors contribute to the NC rate. As a result, NC1 interactions can add a significant background contribution for muon-disappearance samples. In the T2K experiment, presented next in this Letter, statistical and systematic uncertainties for this channel are comparable. In the near future, Hyper-Kamiokande (HK) and DUNE, are expected to select hundreds of NC1 events as part of their oscillation samples. Therefore, better understanding the NC1 cross section and constraining its modeling uncertainty is important for the precision oscillation physics goals of next-generation experiments.
Experimental Setup. The Tokai-to-Kamioka (T2K) long-baseline neutrino experiment Abe et al. (2011) studies neutrino oscillations using a highly pure muon (anti)neutrino beam produced by the neutrino facility at the J-PARC proton accelerator. Data collection is divided in -beam or -beam runs. The experiment combines measurements at two detector sites: the near detector facility, located 280 meters from the proton beam target, including the main near detector ND280, which was used to collect the data for this study; and the far detector facility, located 295 km downstream, is the Super-Kamiokande (SK) detector Fukuda et al. (2003). Both ND280 and SK are placed 2.5∘ off-axis, resulting in a narrow neutrino energy spectrum peaking at 0.6 GeV Abe et al. (2013); Beavis et al. (1995).
The analysis utilizes ND280 data recorded in -beam mode. ND280 Abe et al. (2011) is a magnetized particle detector made of several sub-detector modules, including: two Fine-Grained-Detectors (FGD1 and FGD2) Amaudruz et al. (2012), three gaseous Time-Projection-Chambers (TPCs) Abgrall et al. (2011) and an Electromagnetic Calorimeter (ECal) Allan et al. (2013). Together, they provide rich information used to reconstruct the particle type, its trajectory and its momentum.
Analysis samples. Events with a vertex reconstructed in the fully-active hydrocarbon target FGD1 are classified in four different regions: a signal-enriched sample and three background-enriched samples. The signal definition includes all and interactions without charged leptons in the final state, a single positive pion, any number of protons with momentum below the detection threshold of 200 MeV/c (See details in Ref. Abe et al. (2025)), and no other mesons or other charged particles in the final state. Due to ND280’s acceptance and particle identification capabilities, the measurement is reported in the kinematic Region of Interest (RoI) defined by the conditions 0.5 and GeV/. The neutrino generator NEUT v5.4.0 Hayato (2009a); Hayato and Pickering (2021) is used with T2K’s flux and detector simulation pipeline Abe et al. (2011) to predict the rate of these events. Over 200 signal events are expected overall, with approximately 165 in the signal sample at 30.5% efficiency and 51.4% purity, and an integrated detector uncertainty for the predicted event rate below 5%. Three analysis sidebands enriched in background events are used in the analysis. This includes a sample with 81.0% purity in CC events, the most numerous background events in the signal sample corresponding to 28.2% of its selected events. The other sidebands are enriched in events with additional tracks in the neutrino target and events where the main track is identified as a proton instead of a . Full selection details and performance metrics are presented in Ref. Abe et al. (2025).
Analysis strategy. The NC1 cross section is measured using an unregularized binned maximum likelihood fit, the same method described in earlier T2K measurements, e.g. in Refs. Abe et al. (2023c, b). The procedure consists of numerically optimizing the parameters of interest, i.e. scaling factors for the number of signal events, together with nuisance parameters describing plausible variations of T2K’s flux, detector and background cross-section models. The agreement with data is simultaneously maximized in the signal and background-enriched samples. This method provides a data-driven background and flux constraint while unfolding detector effects and providing a post-fit prediction of the expected signal events in every analysis bin, which is then used to calculate the signal cross section. In total, the flux-averaged cross section is unfolded in 13 bins in pion angle and momentum, nine of them contained in the RoI. The binning scheme is presented in Fig. 2.
The model independence of the cross-section extraction was validated through extensive tests using alternative datasets with controlled variations. In all cases, small or negligible measurement biases were observed, confirming the robustness of the method.
Figure 2: Scheme of the true pion kinematic bins where the cross section is measured. Bin IDs are shown within their associated kinematic regions. The region of interest, were we report the measurement, is enclosed by an orange solid line and corresponds to bins 2-10. Color shades are used to aid visualization but have no physical meaning.
Interaction Models. We compare the extracted double-differential and integrated cross section to several interaction generators using NUISANCE Stowell et al. (2017): NEUT v5.6.2 Hayato (2002, 2009b); Hayato and Pickering (2021), NuWro v19.02.2 Golan et al. (2012a, b) and GENIEv3 Andreopoulos et al. (2010, 2015) using the AR23_20i configuration. These represent a range of commonly used models, which differ in a variety of details including hard scattering processes, nuclear modeling and re-interactions of interaction products within the nucleus (Final State Interactions, FSI). Additionally, we investigate the role of FSI using five different GENIE model variations, where the only change is the FSI model used. These represent a broad range of models of varying levels of sophistication ranging from no FSI, and full semi-classical cascade models. They include: the INTRANUKE Dytman (2009) hA single-step empirical model (G18_10a); the INTRANUKE hN full intranuclear cascade (G18_10b); GENIE’s implementation of the INCL cascade Mancusi et al. (2015); Cugnon et al. (2016) (G18_10c); the GEANT4 Bertini cascade Wright and Kelsey (2015) (G18_10d); and an unphysical model in which all final-state interactions have been turned off (G18_10X). AR23_20i uses the hA model and is identical to G18_10a for NC pion production, except that the free nucleon pion production tuning from Ref. Tena-Vidal et al. (2021) is applied.
Figure 3: Measured NC1 double-differential cross section compared to generator predictions.Figure 4: Measured NC1 double-differential cross section compared to various FSI predictions.
Results. The measured data cross section is compared to different neutrino interaction generator predictions in Fig. 3 and FSI models in Fig. 4. The data agrees well with generator and FSI model predictions, with values below nine, the number of differential bins, in all cases. Bins 5, 6, 8, and 9, which cover the kinematic region with highest purity and signal abundance, all prefer a higher signal cross section than that of all models. The data preference for a larger cross section is emphasized when studying the integrated cross section, summarized in Tab. 1. Among all the models investigated, GENIEv3 AR23_20i and G18_10X show the largest disagreement, both at a statistical tension with data equivalent to 1.7. The best agreement between data and simulation for the integrated cross section is found with G18_10d.
Table 1: Integrated cross section values in the RoI with their associated values for different generators and FSI models.
Model
(cm2 / nucleon)
p-value
Data
GENIEv3 AR23_20i_00_000
0.093
NuWro v19.02.2
0.153
NEUT v5.6.2
0.108
G18_10X
0.089
G18_10a
0.100
G18_10b
0.133
G18_10c
0.176
G18_10d
0.507
Conclusions. Using the largest data sample of selected NC interactions to date, the double-differential cross section for this channel has been presented for the first time. Overall, the differential cross-section results agree with tested model predictions, though we observe a statistically weak preference for a cross section approximately 30% higher than most model expectations. The variations observed across FSI models are greater than those between neutrino interaction generators, and could significantly reduce the observed normalization discrepancy. This highlights the critical need to enhance our understanding of final-state interactions. We anticipate that NC predictions will benefit from forthcoming high-statistics neutrino interaction measurements from T2K and other experiments. In particular, future measurements that precisely measure final-state hadrons in interactions both with and without final-state pions and utilizing transverse kinematic imbalance observables will be especially valuable to inform existing nuclear ground state and final-state-interaction models, directly benefiting NC model predictions.
This work represents the first study of this interaction primarily with sub-GeV neutrinos, enabled by T2K’s neutrino beamline, with dedicated emphasis on achieving a model-independent cross-section extraction, as discussed in Ref. Abe et al. (2025). The results can be used to constrain NC background events for T2K and the future Hyper-Kamiokande experiment, planned to start operations in 2027.
DUNE, using an Argon target and a higher neutrino energy, will likely benefit from additional studies on this channel. The selection strategy developed for this study, may guide future NC measurements at other experiments.
This work establishes the foundation for more precise measurements of this channel in T2K and Hyper-Kamiokande using the ND280 detector. With the recent upgrade of the ND280 detector Abe et al. (2019) and planned increases in data collection over the coming years, we expect continued improvements in the understanding of this channel, enabling more precise measurements of neutrino oscillation parameters. We anticipate that the upgraded ND280 geometry, with an enhanced angular acceptance, should allow to greatly extend the phase space of this measurement in the future to regions where the pion is emitted at large angles with respect to the beam, perhaps even including samples with backwards-going pions. An increase in the signal purity is also expected thanks to the better granularity and timing capabilities of the SuperFGD hydrocarbon target and the addition of Time-of-flight panels in the detector. Unlike the target used in this analysis, SuperFGD is also capable of identifying neutrons, opening new avenues to further understand this channel.
The measurement results have an associated data release that can be found in Ref. T2K Collaboration (2025).
Acknowledgments
The T2K collaboration would like to thank the J-PARC staff for superb accelerator performance. We thank the CERN NA61/SHINE Collaboration for providing valuable particle production data. We acknowledge the support of MEXT, JSPS KAKENHI and bilateral programs, Japan; NSERC, the NRC, and CFI, Canada; the CEA and CNRS/IN2P3, France; the Deutsche Forschungsgemeinschaft (DFG 397763730, 517206441), Germany; the NKFIH (NKFIH 137812 and TKP2021-NKTA-64), Hungary; the INFN, Italy; the Ministry of Science and Higher Education (2023/WK/04) and the National Science Centre (UMO-2018/30/E/ST2/00441, UMO-2022/46/E/ST2/00336 and UMO-2021/43/D/ST2/01504), Poland; the RSF (RSF 24-12-00271) and the Ministry of Science and Higher Education, Russia; MICINN, ERDF and European Union NextGenerationEU funds and CERCA program and Generalitat de Catalunya, Spain; the SNSF and SERI, Switzerland; the STFC and UKRI, UK; the DOE, USA; and NAFOSTED (103.99-2023.144,IZVSZ2.203433), Vietnam. We also thank CERN for the UA1/NOMAD magnet, DESY for the HERA-B magnet mover system, the BC DRI Group, Prairie DRI Group, ACENET, SciNet, and CalculQuebec consortia in the Digital Research Alliance of Canada, and GridPP in the United Kingdom, the CNRS/IN2P3 Computing Center in France and NERSC, USA. In addition, the participation of individual researchers and institutions has been further supported by funds from the ERC (FP7), “la Caixa” Foundation, the European Union’s Horizon 2020 Research and Innovation Programme under the Marie Sklodowska-Curie grant; the JSPS, Japan; the Royal Society, UK; French ANR and Sorbonne Université Emergences programmes; the VAST-JSPS (No. QTJP01.02/20-22); and the DOE Early Career programme, USA. For the purposes of open access, the authors have applied a Creative Commons Attribution license to any Author Accepted Manuscript version arising.
Esteban et al. (2024)I. Esteban, M. C. Gonzalez-Garcia, M. Maltoni, I. Martinez-Soler, J. a. P. Pinheiro, and T. Schwetz, (2024), arXiv:2410.05380 [hep-ph] .
Andreopoulos et al. (2019)C. Andreopoulos et al. (NuSTEC), in NuSTEC Workshop on Shallow- and Deep-Inelastic Scattering (2019) arXiv:1907.13252 [hep-ph] .
Aliaga et al. (2020)L. Aliaga et al. (NuSTEC), in NuSTEC Workshop on Neutrino-Nucleus Pion Production in the Resonance Region (2020) arXiv:2011.07166 [hep-ph] .
Andreopoulos et al. (2015)C. Andreopoulos, C. Barry, S. Dytman, H. Gallagher, T. Golan, R. Hatcher, G. Perdue, and J. Yarba, (2015), arXiv:1510.05494 [hep-ph] .
Dytman (2009)S. Dytman, Neutrino interactions: From theory to Monte Carlo simulations. Proceedings, 45th Karpacz Winter School in Theoretical Physics, Ladek-Zdroj, Poland, February 2-11, 2009, Acta Phys. Polon. B40, 2445 (2009).
Mancusi et al. (2015)D. Mancusi, A. Boudard, J. Carbonell, J. Cugnon, J.-C. David, and S. Leray, Phys. Rev. C 91, 034602 (2015).